OpenStudy (anonymous):

Write the equation of the line in Slope-Intercept form, Then graph the equation by using either the slope-intercept equation or by using x- and y-intercepts. The line contains points (5, 6) and (2, 6).

OpenStudy (anonymous):

First you need to find the slope

OpenStudy (anonymous):

It is, 3 I think....

OpenStudy (anonymous):

(6-6)/(2-5)=?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

It would be 0.

OpenStudy (anonymous):

Well I feel stupid.

OpenStudy (anonymous):

It's alright. Now that you have the slope, you can use that plus one of your points to get slope intercept form. use: y-y1=m(x-x1) The (x1, y1) is one of your points and m is the slope. So plug those in

OpenStudy (anonymous):

1-y1=m(1-x1) ?

OpenStudy (anonymous):

Noo, m=slope and (x1, y1) is whichever point you wanna use. So well use (5, 6) y-6=0(x-5)

OpenStudy (anonymous):

When you distribute the 0 on the right it all becomes 0, then add the 6 and you get: y=6

OpenStudy (anonymous):

Ahh, okay. I see!

OpenStudy (anonymous):

Alright, and there was kinda a shortcut to this one. You see on both the points your y coordinate is 6. Since its a line that means every y coordinate will be 6. So, the equation is y=6 Dunno if that helped or confused you. But im glad you understand the first part

OpenStudy (anonymous):

Yeah, I noticed that. This one seemed to be pretty simple actually.